REINVESTIGATION OF THE MICROWAVE SPECTRUM OF ACETAMIDE

R. D. SUENRAM, \underline{G. Yu. GOLUBIATNIKOV}, I. I. LEONOV, J. T. HOUGEN, G. T. FRASER, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441; I. KLEINER, Laboratoire de Photophysique Moléculaire, Université Paris-Sud, 91405 Orsay, France; J. ORTIGOSO, Instituto de Estructura de la Materia, CSIC, Serrano 121, 28006 Madrid, Spain.

About 50 jet-cooled Fourier transform lines for acetamide have been recorded using a new version of our spectrometer, which has been upgraded with a heated nozzle and an expanded automatic scanning range. Nuclear quadrupole hyperfine structure arising from the nitrogen atom was removed theoretically to yield hyperfine-free center frequencies. In addition, about 30 submillimeter measurements were carried out. When hyperfine structure was observed for these lines, it was also removed theoretically. A set of 115 A-species and E-species rotational transitions in the torsional ground state, obtained by combining our new measurements with the literature data, have been fit to a model involving 28 torsion, rotation, and torsion-rotation interaction parameters to near experimental uncertainty (i.e., to a weighted unitless standard deviation of 1.5), significantly improving on previous fits. Various theoretical problems associated with K labels for E-species levels arise in this very low barrier molecule, requiring a variant of the signed K_{a} labels frequently used for internal-rotor E states.